# Quillen’s conjecture and unitary groups

### v3

@@ 2.4. Symmetric group and certain subsets of it
Thus by the inductiom hypothesis for n1n-1 applied to $w^{\prime}=w_{2}\cdots w_{n-2}\in S_{n-1}^{-}$, we have that $w^{\prime}(\{j,j+1\})=\{1,2\}$, and then $w(\{j,j+1\})=w^{\prime}(\{j,j+1\})=\{1,2\}$. If j=k1j=k-1, then
@@
$s_{1}w_{2}\cdots w_{n-2}w_{n-1}s_{k}$ $=s_{1}w_{2}\cdots w_{n-2}s_{n-1}\cdots s_{k}s_{j}$
@@
Very roughly, a simplicial complex is a type of topological space made out of vertices, segments, triangles, tetrahedrons, etc., as well as their $n$-dimensional counterparts, which are called $n$-simplices. We regard the topological realization $|\mathcal{A}_{p}(G)|$ as a simplicial complex, so that it has one ordered $n$-simplex for each chain of length $n$ in $\mathcal{A}_{p}(G)$,
@@
We note that the faces of such an $n$-simplex are the ordered (n1)(n-1)-simplices
@@
$d_{i}(P_{0}<\ldots<P_{n})=P_{0}<\cdots<\widehat{P_{i}}<\cdots<P_{n},$
@@
Its differential $d(\mathbf{C}_{E})\in C_{r-2}(|\mathcal{A}_{p}(E)|;R)$ is described in the next result, cf. [6, Proposition 3.2], [7, Proposition 3.2].
@@
$C_{x,\mathbf{i}}$ $=\sum_{(x,\mathbf{j})\in\mathcal{C}(x,\mathbf{i})}h(x)\operatorname{sgn}(\mathbf{j})\text{, }\quad\text{and}$
$D_{x,\mathbf{i}}$ $=(-1)^{r-1}\sum_{(x,\mathbf{j})\in\mathcal{D}(x,\mathbf{i})}h(x)\operatorname{sgn}(\mathbf{j}),$
@@
$\mathcal{C}(x,\mathbf{i})$ $=\{(x,\mathbf{j})\in X\times\mathcal{T}^{r}~|~{}^{{x}\!}{\sigma}_{\mathbf{j}}={}^{{x}\!}{\sigma}_{\mathbf{i}}\}\quad\hbox{and}\quad$
$\mathcal{D}(x,\mathbf{i})$ $=\{(x,\mathbf{j})\in X\times\mathcal{T}^{r}~|~{}^{{x}\!}{\tau}_{\mathbf{j}}={}^{{x}\!}{\tau}_{\mathbf{i}}\}.$
If we further assume that $E$ is a maximal elementary abelian $p$-subgroup of $G$, we immediately obtain the following result, see also [7, Theorem 3.3].
@@
    $C_{x,\mathbf{i}}\neq 0$ for some $x\in X$ and some $\mathbf{i}\in\mathcal{T}^{r}$,
@@
    $D_{x,\mathbf{i}}=0$ for all $x\in X$ and all $\mathbf{i}\in\mathcal{T}^{r}$.
@@
The maximality condition on $E$ is needed to ensure that the non-trivial cycle $\mathbf{C}_{E,X,h}]$ is not a boundary. This maximality condition holds, for instance, if $E$ attains the maximum possible rank among the elementary $p$-subgroups of $G$, i.e., if $r=\operatorname{m}_{p}(G)$. The conclusion of the theorem implies that $|\mathcal{A}_{p}(G)|$ is not contractible and, in particular, that Quillen’s conjecture holds for $G$. If $r=\operatorname{m}_{p}(G)$ and $R=\mathbb{Z}$ or $\mathbb{Q}$, the theorem shows that the $\mathcal{Q}\mathcal{D}_{p}$ property holds, see [3, Definition p.474].
@@
The finite group $G$ with $r=\operatorname{m}_{p}(G)$ has the Quillen dimension at $p$ property, $\mathcal{Q}\mathcal{D}_{p}$ for short, if
@@
The case $R=\mathbb{Z}$ implies $\mathcal{Q}\mathcal{D}_{p}$ as top dimension integral homology is a free abelian group. Note also that the hypotheses of Theorem 3.9 imply that $O_{p}(G)=1$ because of Quillen’s “conical contractibility” [19, Proposition 2.4]. Next, we describe a set of general group theoretical conditions that imply item (a) or (b) of Theorem 3.9. Point (b2) below will be used in the applications in Section 4.
@@
Let $E=\langle e_{1},\dots,e_{r}\rangle$ be a maximal elementary abelian $p$-subgroup of rank $r$ of the group $G$, $\{x_{1},\ldots,x_{r}\}\subseteq X\subseteq G$, and $h\colon X\to R^{*}$. Then:
@@
    If for all $1\leq i\leq r$, $x_{i}\in C_{G}(E_{i})$, and, for all such $i$ and all $\in X, xx$$x_{i}^{\epsilon}\in X$ for exactly one value $\epsilon\in\{-1,+1\}$ and $h(x)+h(xx_{i}^{\epsilon})=0$, then $G$, $E$, $X$, $h$ satisfy hypothesis (b) of Theorem 3.9.
@@
    If for all for $1\leq i\leq r$, $x_{i}\in C_{G}(E_{i})$ and $x_{i}$ has order $2$, and, for all such $i$ and all $\in X, xx$$x_{i}\in X$ and $h(x)+h(xx_{i})=0$, then $G$, $E$, $X$, $h$ satisfy hypothesis (b) of Theorem 3.9.
@@
We consider the chain $\mathbf{C}_{E,X,h}$ of Definition 3.8 and start with item (a): Consider $x\in X$ and $\mathbf{i}\in\mathcal{T}^{r}$. If $(x^{\prime},\mathbf{j})\in\mathcal{C}(x,\mathbf{i})$ then $^{{x^{\prime}}\!}{\sigma}_{\mathbf{j}}={}^{{x}\!}{\sigma}_{\mathbf{i}}$ and, in particular, $^{{x^{\prime}}\!}{E}={}^{{x}\!}{E}$. Hence $x^{-1}x^{\prime}\in N_{G}(E)$ and, by hypothesis (a) in the statement, we must have $x^{\prime}=x$. Thus, we have $\mathbf{i}=\mathbf{j}$, $\mathcal{C}(x,\mathbf{i})=\{(x,\mathbf{i})\}$, and $C(x,\mathbf{i})=h(x)\neq 0$.
For item (b1), we decompose $\mathcal{D}(x,\mathbf{i})=\cup_{\mathbf{j}\in\mathcal{T}^{r}}\mathcal{D}(x,\mathbf{i},\mathbf{j})\times\{\mathbf{j}\}$ with
$\mathcal{D}(x,\mathbf{i},\mathbf{j})=\{y\in X~|~{}^{{y}\!}{\tau}_{\mathbf{j}}={}^{{x}\!}{\tau}_{\mathbf{i}}\}.$
Then $D(x,\mathbf{i})=(-1)^{r-1}\sum_{\mathbf{j}\in\mathcal{T}^{r}}\operatorname{sgn}(\mathbf{j})D(x,\mathbf{i},\mathbf{j})$ with $D(x,\mathbf{i},\mathbf{j})=\sum_{y\in\mathcal{D}(x,\mathbf{i},\mathbf{j})}h(y)$. Hence, to check that $D(x,\mathbf{i})=0$, it is enough to check that $D(x,\mathbf{i},\mathbf{j})=0$ for all $\mathbf{j}$.
So let $\mathbf{j}=[j_{1},\ldots,j_{r-1}]\in\mathcal{T}^{r}$. Then, by hypothesis (b1), the element $x_{j_{1}}$ belongs to $C_{G}(E_{j_{1}})=C_{G}(\tau_{\mathbf{j}})$. For any element $y\in\mathcal{D}(x,\mathbf{i},\mathbf{j})$ we have, by (b1) again, that $yx_{j_{1}}^{\epsilon}\in X$ for $\epsilon$ equal to $1$ or 1-1. As $x_{j_{1}}\in C_{G}(\tau_{\mathbf{j}})$, we also have that $yx_{j_{1}}^{\epsilon}\in\mathcal{D}(x,\mathbf{i},\mathbf{j})$. So, if we define,
$\mathcal{D}^{+}(x,\mathbf{i},\mathbf{j})=$ $\{y\in\mathcal{D}(x,\mathbf{i},\mathbf{j})~|~yx_{j_{1}}\in\mathcal{D}(x,\mathbf{i},\mathbf{j})\}\text{ and }$
$\mathcal{D}^{-}(x,\mathbf{i},\mathbf{j})=$ $\{y\in\mathcal{D}(x,\mathbf{i},\mathbf{j})~|~yx_{j_{1}}^{-1}\in\mathcal{D}(x,\mathbf{i},\mathbf{j})\},$
we have $\mathcal{D}(x,\mathbf{i},\mathbf{j})=\mathcal{D}^{+}(x,\mathbf{i},\mathbf{j})\cup\mathcal{D}^{-}(x,\mathbf{i},\mathbf{j})$. Also, if $y\in\mathcal{D}^{+}(x,\mathbf{i},\mathbf{j})\cap\mathcal{D}^{-}(x,\mathbf{i},\mathbf{j})$, then both $yx_{j_{1}}$ and $yx_{j_{1}}^{-1}$ belong to $X$, contradicting the uniqueness of $\epsilon$ in point (b1). So this intersection is empty. In addition, the maps $\mathcal{D}^{+}(x,\mathbf{i},\mathbf{j})\to\mathcal{D}^{-}(x,\mathbf{i},\mathbf{j})$, $y\mapsto yx_{j_{1}}$, and $\mathcal{D}^{-}(x,\mathbf{i},\mathbf{j})\to\mathcal{D}^{+}(x,\mathbf{i},\mathbf{j})$, $y\mapsto yx_{j_{1}}^{-1}$, are inverse to each other and hence bijections. Therefore, $\mathcal{D}(x,\mathbf{i},\mathbf{j})=\cup_{y\in\mathcal{D}^{+}(x,\mathbf{i},\mathbf{j})}\{y,yx_{i_{1}}\}$ and
$D(x,\mathbf{i},\mathbf{j})=\sum_{y\in\mathcal{D}^{+}(x,\mathbf{i},\mathbf{j})}h(y)+h(yx_{j_{1}})=0,$
where we have used item (b1) once more. Finally, for point (b2), the arguments are identical but for considering the non-trivial involution $\mathcal{D}(x,\mathbf{i},\mathbf{j})\to\mathcal{D}(x,\mathbf{i},\mathbf{j})$, $y\mapsto yx_{j_{1}}$, and the corresponding partition into subsets of the form $\{y,yx_{j_{1}}\}$. ∎
Note that, if $X$ is a subgroup, the condition of Lemma 3.11(a) takes the following simplified form, and an example with such a subgroup $X$ is given below,
$X\cap N_{G}(E)=\{1\}.$
@@
Then it is straightforward that items (a) and (b2) in Lemma 3.11 hold, so that Theorem 3.9 gives $\widetilde{H}_{1}(|\mathcal{A}_{3}(G)|;\mathbb{Z})\neq 0$ and that $\mathcal{Q}\mathcal{D}_{3}$ holds for $G$. In Figure 3, the non-zero homology cycle constructed is depicted, with bullet points and squares corresponding, respectively, to $p$-rank 1 and 2 elementary abelian $p$-subgroups. See also the barycentric subdivisions in Remark 3.7.
@@
Figure 3. Triangulation of the sphere $S^{1}$ arising from $S_{7}$.
As explained in [7, Remark 4.3], in this case we cannot expect to find elements $x_{1}$ and $x_{2}$ that commute.
@@
In this section, we prove the Quillen dimension property for $\operatorname{PGU}\nolimits_{n}(q)$ and its extensions by field automorphisms of order $p$, as in Theorem A of the Introduction. The techniques we employ are better described in terms of the symmetric group embedded into $\operatorname{GU}\nolimits_{n}(q)$, see Definition 4.1, and quasi-reflections, see Subsection 2.3. We will apply Theorem 3.9 and Lemma 3.11 to construct non-trivial homology classes in the appropriate dimensions; see Theorem 4.3 for the case of $\operatorname{PGU}\nolimits_{n}(q)$ and Theorem 4.7 for the case of its extension by field automorphisms of order $p$. In general, there will be non-trivial identifications among simplices. See Remark 4.8 and Example 4.9 for more details. Unless stated otherwise, we assume throughout this section that $p$ is an odd prime dividing $q+1$ with $(p,q)\neq(3,2)$, and that $n\geq 2$.
@@
We write $Z=Z(\operatorname{GU}\nolimits_{n}(q))$ and, for an element $x\in\operatorname{GU}\nolimits_{n}(q)$, a subset $X\subseteq\operatorname{GU}\nolimits_{n}(q)$, a function $h\colon X\to\mathbb{Z}^{*}$, or an $r$-simplex of $|\mathcal{A}_{p}(\operatorname{GU}\nolimits_{n}(q))|$
@@
we denote by $[x]$ (resp. $[X]$) the image of $x$ (resp. $X$) in the quotient $\operatorname{PGU}\nolimits_{n}(q)$, we denote by $h]\colon[X]\to\mathbb{Z}^{*}$ the map $[h]([x])=h(x)$, where we assume that $X\to[X]$ is an injection whenever we consider $[h]$, and we denote by $\sigma]$ the $r$-simplex of $|\mathcal{A}_{p}(\operatorname{PGU}\nolimits_{n}(q))|$ given as follows,
@@
For $i=1,\ldots,n-1$, let $y_{i}$ be the usual permutation matrix of $\operatorname{GU}\nolimits_{n}(q)$ that corresponds to the transposition $s_{i}=(i,i+1)\in S_{n}$, and set $Y=\langle y_{1},\ldots,y_{n-1}\rangle$, so that we have an isomorphism $Y\to S_{n}$ given by $y_{i}\mapsto s_{i}$.
In the rest of the paper, we will employ the isomorphism $Y\to S_{n}$ above without further notice. In particular, we define the subsets $Y^{+},Y^{-},\partial Y^{+},\partial Y^{-}$ as the pre-images of the subsets $S_{n}^{+},S_{n}^{-},\partial S_{n}^{+},\partial S^{-}_{n}$ by this isomorphism, respectively, see Definition 2.6. We next recall that $Y$ acts on diagonal elements of $T\leq\operatorname{GU}\nolimits_{n}(q)$ as follows,
@@
Moreover, there exists a non-diagonal element $x\in\operatorname{GU}\nolimits_{n}(q)$ such that, if we set $X=\{x^{\delta}y\mid\delta=0,1,y\in Y^{-}\}$, we have,
@@
Next, we define the element $x$ of the statement as the quasi-reflection given by Proposition 2.3, so that point (d) is immediate from the conclusions there. For point (c), assume that
$x^{\delta}y=zx^{\delta^{\prime}}y^{\prime}\in Z$
for $\delta,\delta^{\prime}=0,1$, $y,y^{\prime}\in Y^{-}$, and $z\in Z$. Then $x^{\delta-\delta^{\prime}}=zy^{\prime}y^{-1}$ and, by point (d), the action by conjugation of the left-hand side fixes every element of the form $t=\operatorname{diag}(\alpha_{1},\alpha_{1},\alpha_{3},\ldots,\alpha_{n})\in T$, so that, by Equation (12), $y^{\prime}=y$ or $y^{\prime}=y_{1}y$. This latter equality implies that $y^{\prime}\in Y^{+}$ by the third part of Proposition 2.8(a), and this cannot hold because of the second part of Proposition 2.8(a). Thus $x^{\delta-\delta^{\prime}}=z$ is diagonal, which implies that $\delta=\delta^{\prime}$, as $x$, and hence also $x^{-1}$, are not diagonal. ∎
@@
If $n\geq 2$, $p$ is odd, $p$ divides $q+1$, and $(p,q)\neq(3,2)$, then we have $H_{n-2}(|\mathcal{A}_{p}(\operatorname{PGU}\nolimits_{n}(q))|;\mathbb{Z})\neq 0$ and $\mathcal{Q}\mathcal{D}_{p}$ holds for $\operatorname{PGU}\nolimits_{n}(q)$.

### v5

@@ 2.4. Symmetric group and certain subsets of it
Thus, by the inductiom hypothesis for n1n-1 applied to $w^{\prime}=w_{2}\cdots w_{n-2}\in S_{n-1}^{-}$, we have that $w^{\prime}(\{j,j+1\})=\{1,2\}$, and then $w(\{j,j+1\})=w^{\prime}(\{j,j+1\})=\{1,2\}$. If j=k1j=k-1, then
@@
$s_{1}w_{2}\cdots w_{n-2}w_{n-1}s_{j}$ $=s_{1}w_{2}\cdots w_{n-2}s_{n-1}\cdots s_{k}s_{j}$
@@
We regard the topological realization $|\mathcal{A}_{p}(G)|$ as a simplicial complex, so that it has one ordered $n$-simplex for each chain of length $n$ in $\mathcal{A}_{p}(G)$,
@@
We note that the maximal faces of such an $n$-simplex are the ordered (n1)(n-1)-simplices
@@
$d_{i}(P_{0}<\cdots<P_{n})=P_{0}<\cdots<\widehat{P_{i}}<\cdots<P_{n},$
@@
Its differential $d(\mathbf{C}_{E})\in C_{r-2}(|\mathcal{A}_{p}(E)|;R)$ is described in the next result, cf. [7, Proposition 3.2], [8, Proposition 3.2].
@@
$C(x,\mathbf{i})$ $=\sum_{(y,\mathbf{j})\in\mathcal{C}(x,\mathbf{i})}h(y)\operatorname{sgn}(\mathbf{j})\text{, }\quad\text{and}$
$D(x,\mathbf{i})$ $=(-1)^{r-1}\sum_{(y,\mathbf{j})\in\mathcal{D}(x,\mathbf{i})}h(y)\operatorname{sgn}(\mathbf{j}),$
@@
$\mathcal{C}(x,\mathbf{i})$ $=\{(y,\mathbf{j})\in X\times\mathcal{T}^{r}~|~{}^{{y}\!}{\sigma}_{\mathbf{j}}={}^{{x}\!}{\sigma}_{\mathbf{i}}\}\quad\hbox{and}\quad$
$\mathcal{D}(x,\mathbf{i})$ $=\{(y,\mathbf{j})\in X\times\mathcal{T}^{r}~|~{}^{{y}\!}{\tau}_{\mathbf{j}}={}^{{x}\!}{\tau}_{\mathbf{i}}\}.$
If we further assume that $E$ is a maximal elementary abelian $p$-subgroup of $G$, we immediately obtain the following result, see also [8, Theorem 3.3].
@@
    $C(x,\mathbf{i})\neq 0$ for some $x\in X$ and some $\mathbf{i}\in\mathcal{T}^{r}$,
@@
    $D(x,\mathbf{i})=0$ for all $x\in X$ and all $\mathbf{i}\in\mathcal{T}^{r}$.
@@
The maximality condition on $E$ is needed to ensure that the non-trivial cycle $\mathbf{C}_{E,X,h}]$ is not a boundary. This maximality condition holds, for instance, if $E$ attains the maximum possible rank among the elementary $p$-subgroups of $G$, i.e., if $r=\operatorname{m}_{p}(G)$. The conclusion of the theorem implies that $|\mathcal{A}_{p}(G)|$ is not contractible and, in particular, that Quillen’s conjecture holds for $G$. If $r=\operatorname{m}_{p}(G)$, then the conclusion of the theorem and the fact that top dimension integral homology is a free abelian group, show that the $\mathcal{Q}\mathcal{D}_{p}$ property holds, see [4, Definition p.474].
@@
The finite group $G$ with $r=\operatorname{m}_{p}(G)$ has the Quillen dimension property for $p$, $\mathcal{Q}\mathcal{D}_{p}$ for short, if
@@
Note also that the hypotheses of Theorem 3.9 imply that $O_{p}(G)=1$ because of Quillen’s “conical contractibility” [19, Proposition 2.4]. Next, we describe a set of general group theoretical conditions that imply item (a) or (b) of Theorem 3.9, and some of the upcoming arguments are based on the following elementary fact.
@@
Let $\mathcal{D}$ be a finite set, $h\colon\mathcal{D}\to R$ a function, and $\Omega$ a non-trivial involution of $\mathcal{D}$, i.e., a map $\Omega\colon\mathcal{D}\to\mathcal{D}$ such that $\Omega^{2}(y)=y\neq\Omega(y)$ for all $y\in\mathcal{D}$, and assume that $h(y)+h(\Omega(y))=0$ for all $y\in\mathcal{D}$. Then $\sum_{y\in\mathcal{D}}h(y)=0$.
Proof.
Partition the set $\mathcal{D}$ into the subsets $\{y,\Omega(y)\}$. ∎
In the statement below, for a subset $X$ and a element $x$ of a group, we write $Xx$ for the subset $\{yx\mid y\in X\}$, and we set, for $\mathbf{i},\mathbf{j}\in\mathcal{T}^{r}$ with $\mathbf{j}=[j_{1},\ldots,j_{r-1}]$,
(10) $\mathcal{D}(x,\mathbf{i},\mathbf{j})$ $=\{y\in X~|~{}^{{y}\!}{\tau}_{\mathbf{j}}={}^{{x}\!}{\tau}_{\mathbf{i}}\}\text{, }\mathcal{D}^{*}(x,\mathbf{i},\mathbf{j})=\mathcal{D}(x,\mathbf{i},\mathbf{j})\cap(Xx_{j_{1}}\cup Xx_{j_{1}}^{-1}),$
[…]
Lemma 3.11, Equation (10), and items (b,b2,c) below will be used in Section 4. We stress that, in the statement below, the elements $x_{i}$ do not have to belong to $X$.
Lemma 3.12.
Let $E=\langle e_{1},\dots,e_{r}\rangle$ be a maximal elementary abelian $p$-subgroup of rank $r$ of the group $G$, $\{x_{1},\ldots,x_{r}\}\subseteq G$, $X\subseteq G$, and $h\colon X\to R^{*}$. Then
@@
Moreover, if $\mathbf{i},\mathbf{j}\in\mathcal{T}^{r}$, $\mathbf{j}=[j_{1},\ldots,j_{r-1}]$, and we assume that
  1. (b)
    $x_{j_{1}}\in C_{G}(E_{j_{1}})$ and, if $y\in X\cap Xx_{j_{1}}^{\epsilon}$, $\epsilon\in\{-1,1\}$, then $h(y)+h(yx_{j_{1}}^{-\epsilon})=0$,
then the following hold,
@@
    If $X\cap Xx_{j_{1}}\cap Xx_{j_{1}}^{-1}=\emptyset$, then $D^{*}(x,\mathbf{i},\mathbf{j})=0$.
@@
    If $x_{j_{1}}$ has order $2$, then $D^{*}(x,\mathbf{i},\mathbf{j})=0$.
In addition, the following statements hold,
  1. (c)
    $D(x,\mathbf{i})=(-1)^{r-1}\sum_{\mathbf{j}\in\mathcal{T}^{r}}\operatorname{sgn}(\mathbf{j})D(x,\mathbf{i},\mathbf{j})$.
  2. (c1)
    If for all $1\leq j_{1}\leq r$, $X\subseteq(Xx_{j_{1}}\cup Xx_{j_{1}}^{-1})$ and the hypotheses of (b) and (b1) hold, then $G$, $E$, $X$, $h$ satisfy hypothesis (b) of Theorem 3.9.
  3. (c2)
    If for all $1\leq j_{1}\leq r$, $X\subseteq Xx_{j_{1}}$ and the hypotheses of (b) and (b2) hold, then $G$, $E$, $X$, $h$ satisfy hypothesis (b) of Theorem 3.9.
@@
We consider the chain $\mathbf{C}_{E,X,h}$ of Definition 3.8 and start with item (a): Consider $x\in X$ and $\mathbf{i}\in\mathcal{T}^{r}$. If $(x^{\prime},\mathbf{j})\in\mathcal{C}(x,\mathbf{i})$ then $^{{x^{\prime}}\!}{\sigma}_{\mathbf{j}}={}^{{x}\!}{\sigma}_{\mathbf{i}}$ and, in particular, $^{{x^{\prime}}\!}{E}={}^{{x}\!}{E}$. Hence $x^{-1}x^{\prime}\in N_{G}(E)$ and, by the hypothesis in (a), we must have $x^{\prime}=x$. Thus, we have $\mathbf{i}=\mathbf{j}$, $\mathcal{C}(x,\mathbf{i})=\{(x,\mathbf{i})\}$, and $C(x,\mathbf{i})=h(x)\neq 0$.
For item (b1), note that, by definition, $\mathcal{D}^{*}(x,\mathbf{i},\mathbf{j})=\mathcal{D}^{+}(x,\mathbf{i},\mathbf{j})\cup\mathcal{D}^{-}(x,\mathbf{i},\mathbf{j})$, where
$\mathcal{D}^{+}(x,\mathbf{i},\mathbf{j})=\mathcal{D}^{*}(x,\mathbf{i},\mathbf{j})\cap Xx_{j_{1}}\text{ and }\mathcal{D}^{-}(x,\mathbf{i},\mathbf{j})=\mathcal{D}^{*}(x,\mathbf{i},\mathbf{j})\cap Xx_{j_{1}}^{-1},$
and that, by the hypothesis in point (b1),
$\mathcal{D}^{+}(x,\mathbf{i},\mathbf{j})\cap\mathcal{D}^{-}(x,\mathbf{i},\mathbf{j})\subseteq X\cap Xx_{j_{1}}\cap Xx_{j_{1}}^{-1}=\emptyset.$
Now, by the first hypothesis in point (b), $x_{j_{1}}$ belongs to $C_{G}(E_{j_{1}})=C_{G}(\tau_{\mathbf{j}})$. In particular, we have a non-trivial involution $\mathcal{D}^{*}(x,\mathbf{i},\mathbf{j})\to\mathcal{D}^{*}(x,\mathbf{i},\mathbf{j})$, given by $y\mapsto yx_{j_{1}}^{-1}$ if $y\in\mathcal{D}^{+}(x,\mathbf{i},\mathbf{j})$ and by $y\mapsto yx_{j_{1}}$ if $y\in\mathcal{D}^{-}(x,\mathbf{i},\mathbf{j})$. Then we are done by Lemma 3.11 and by the second hypothesis in point (b). For item (b2), the arguments are identical but for considering the non-trivial involution $\mathcal{D}^{*}(x,\mathbf{i},\mathbf{j})\to\mathcal{D}^{*}(x,\mathbf{i},\mathbf{j})$ given by $y\mapsto yx_{j_{1}}$.
Item (c) follows from the decomposition $\mathcal{D}(x,\mathbf{i})=\cup_{\mathbf{j}\in\mathcal{T}^{r}}\mathcal{D}(x,\mathbf{i},\mathbf{j})\times\{\mathbf{j}\}$. Finally, for items (c1) and (c2), note that, in these cases, we have $\mathcal{D}^{*}(x,\mathbf{i},\mathbf{j})=\mathcal{D}(x,\mathbf{i},\mathbf{j})$ and $D^{*}(x,\mathbf{i},\mathbf{j})=D(x,\mathbf{i},\mathbf{j})$, so that the conclusions follow directly from items (b,b1,c) or (b,b2,c), respectively, ∎
Note that, if $X$ is a non-trivial subgroup of $G$ with $\{x_{1},\ldots,x_{r}\}\subseteq X$, then the condition of Lemma 3.12(a) takes the following simplified form,
$X\cap N_{G}(E)=\{1\},$
the hypothesis in Lemma 3.12(b1) does not hold, and the conditions involving $X$ in Lemma 3.12(c1,c2) are valid. Below we give an example with $X$ a subgroup.
@@
Then it is straightforward that the hypotheses in items (a,b,b2,c2) of Lemma 3.12 hold, so that Theorem 3.9 gives $\widetilde{H}_{1}(|\mathcal{A}_{3}(G)|;\mathbb{Z})\neq 0$ and that $\mathcal{Q}\mathcal{D}_{3}$ holds for $G$. In Figure 3, the non-zero homology cycle constructed is depicted, with bullet points and squares corresponding, respectively, to $p$-rank 1 and 2 elementary abelian $p$-subgroups. See also the barycentric subdivisions in Remark 3.7.
@@
Figure 3. Triangulation of the sphere $S^{1}$ arising from $\mathcal{A}_{3}(S_{7})$.
As explained in [8, Remark 4.3], in this case we cannot expect to find elements $x_{1}$ and $x_{2}$ that commute.
@@
In this section, we prove the Quillen dimension property for $\operatorname{PGU}\nolimits_{n}(q)$ and its extensions by field automorphisms of order $p$, as in Theorem A of the Introduction. The techniques we employ are better described in terms of the symmetric group embedded into $\operatorname{GU}\nolimits_{n}(q)$, see Definition 4.1, and quasi-reflections, see Subsection 2.3. We will apply Theorem 3.9 and Lemmas 3.11 and 3.12 to construct non-trivial homology classes in the appropriate dimensions; see Theorem 4.3 for the case of $\operatorname{PGU}\nolimits_{n}(q)$ and Theorem 4.7 for the case of its extension by field automorphisms of order $p$. In general, there will be non-trivial identifications among simplices. See Remark 4.8 and Example 4.9 for more details. Unless stated otherwise, we assume throughout this section that $p$ is an odd prime dividing $q+1$ with $(p,q)\neq(3,2)$, and that $n\geq 2$.
@@
Throughout this section, we write $Z=Z(\operatorname{GU}\nolimits_{n}(q))$ and, for an element $x\in\operatorname{GU}\nolimits_{n}(q)$ (a subset $X\subseteq\operatorname{GU}\nolimits_{n}(q)$), we denote by $[x]$ ($[X]$) the image of $x$ ($X$) in the quotient $\operatorname{PGU}\nolimits_{n}(q)$. Moreover, for a function $h\colon X\to\mathbb{Z}^{*}$, we denote by $h]\colon[X]\to\mathbb{Z}^{*}$ the map $[h]([x])=h(x)$, where we assume that $X\to[X]$ is an injection whenever we consider $[h]$. Finally, for an $r$-simplex of $|\mathcal{A}_{p}(\operatorname{GU}\nolimits_{n}(q))|$,
@@
we denote by $\sigma]$ the $r$-simplex of $|\mathcal{A}_{p}(\operatorname{PGU}\nolimits_{n}(q))|$ given as follows,
@@
For $i=1,\ldots,n-1$, let $y_{i}$ be the usual permutation matrix of $\operatorname{GU}\nolimits_{n}(q)$ that corresponds to the transposition $s_{i}=(i,i+1)\in S_{n}$, and denote by $Y=\langle y_{1},\ldots,y_{n-1}\rangle$ the group of permutation matrices of $\operatorname{GU}\nolimits_{n}(q)$, so that we have an isomorphism $Y\to S_{n}$ given by $y_{i}\mapsto s_{i}$.
In the rest of the paper, we will employ the isomorphism $Y\to S_{n}$ above without further notice. In particular, we define the subsets $Y^{+},Y^{-},\partial Y^{+},\partial Y^{-}$ as the pre-images of the subsets $S_{n}^{+},S_{n}^{-},\partial S_{n}^{+},\partial S^{-}_{n}$ by this isomorphism, respectively, see Definition 2.6. Next, we recall that the permutation matrix $y\in Y$ acts on diagonal elements of $T\leq\operatorname{GU}\nolimits_{n}(q)$ as follows,
@@
Moreover, there exists an element $x\in\operatorname{GU}\nolimits_{n}(q)$ such that, if we set $X=\{x^{\delta}y\mid\delta=0,1,y\in Y^{-}\}$, we have,
@@
Next, we define the element $x$ of the statement as the quasi-reflection given by Proposition 2.3, so that point (d) is immediate from the conclusions there. Regarding point (c), we prove, for later use, the following stronger claim,
(13) If $x^{\delta}y]=[x^{\delta^{\prime}}y^{\prime}]$ with $\delta,\delta^{\prime}\in\{0,1\}$ and $y,y^{\prime}\in Y$, then $\delta=\delta^{\prime}$ and $y=y^{\prime}$.
For assume that $x^{\delta}y=zx^{\delta^{\prime}}y^{\prime}$, where $\delta,\delta^{\prime},y,y^{\prime}$ are as above and $z\in Z$. Then $x^{\delta-\delta^{\prime}}=zy^{\prime}y^{-1}$, and we note that the right-hand side is a monomial matrix. But $x$, and hence also $x^{-1}$, is not a monomial matrix by Proposition 2.3. Thus, we conclude that $\delta=\delta^{\prime}$, that $y=zy^{\prime}$, and that $y^{\prime}=y$. ∎
@@
If $n\geq 2$, $p$ is odd, $p$ divides $q+1$, and $(p,q)\neq(3,2)$, then we have $\widetilde{H}_{n-2}(|\mathcal{A}_{p}(\operatorname{PGU}\nolimits_{n}(q))|;\mathbb{Z})\neq 0$ and $\mathcal{Q}\mathcal{D}_{p}$ holds for $\operatorname{PGU}\nolimits_{n}(q)$.